The decibel rating D| is related to the sound inte
Answer #1

Soluiton

             The Decibel rating   D   is   related to   sound intensity I by   the formula

              D = 10log_{10}\left ( \frac{I}{10^{-16}} \right )

      a) Let D and    d represent the decibel ratings of sounds of intensity

             I     and    i    respectively

            Using   Properties   of    logarithms , find a   simplified   formula   for   the difference

            between the two ratings D   and   d    in   terms   of    I    and    i

            We have ,

                   D = 10log_{10}\left ( \frac{I}{10^{-16}} \right )              and

                   d = 10log_{10}\left ( \frac{i}{10^{-16}} \right )

                Now , D - d

                     D - d = 10log_{10}\left ( \frac{I}{10^{-16}} \right )- 10log_{10}\left ( \frac{i}{10^{-16}} \right )

                   D - d = 10 \left [ log_{10}\left ( \frac{I}{10^{-16}} \right ) - log_{10}\left ( \frac{i}{10^{-16}} \right ) \right ]

   D - d = 10 \left [ log_{10}I - log_{10}10^{-16} - log_{10}i +log_{10}10^{-16}\right ]   

                     D - d = 10 \left [ log_{10}I - log_{10}i \right ]

                     D - d = 10log_{10}\left ( \frac{I}{i} \right )

       

b) If     a     Sound's intensity quadruples , To find how many decibels louder

does the    sound become

            From    the above    Part , We have

D = 10log_{10}\left ( \frac{I}{10^{-16}} \right )   

              Here , We   have    I = 4

              D = 10log_{10}\left ( \frac{4}{10^{-16}} \right )

   D = 10log_{10}4 - 10log_{10}(10^{-16})

               D = 10log_{10}4 +160log_{10}10

              D = 10log_{10}4 +160

                D = 10\times 0.6020+160

                D = 6.020+160

                 D = 166.020